### v1

# Inferring entropy production in many-body systems using nonequilibrium MaxEnt

Abstract
We propose a method for inferring entropy production (EP) in high-dimensional stochastic systems, including many-body systems and non-Markovian systems with long memory. Standard techniques for estimating EP become intractable in such systems due to computational and statistical limitations. We infer trajectory-level EP and lower bounds on average EP by exploiting a nonequilibrium analogue of the Maximum Entropy principle, along with convex duality. Our approach uses only samples of trajectory observables (such as spatiotemporal correlation functions). It does not require reconstruction of high-dimensional probability distributions or rate matrices, nor any special assumptions such as discrete states or multipartite dynamics. It may be used to compute a hierarchical decomposition of EP, reflecting contributions from different kinds of interactions, and it has an intuitive physical interpretation as a thermodynamic uncertainty relation. We demonstrate its numerical performance on a disordered nonequilibrium spin model with 1000 spins and a large neural spike-train dataset.
The central quantity of interest in nonequilibrium thermodynamics is entropy production (EP) [1]. EP characterizes departure from thermodynamic equilibrium in microscopic physical systems, where it quantifies the dissipation of thermodynamic free energy. More generally, EP provides an information-theoretic measure of temporal irreversibility, including in meso- and macroscopic systems [2, 3, 4, 5]. There is great interest in quantifying EP in high-dimensional complex systems with many degrees of freedom and/or long non-Markovian memory, such as nonequilibrium disordered networks [6], biological active matter [7, 8], and brain networks [9, 10, 11, 12, 13], but this is computationally and statistically difficult. Standard methods require reconstruction of trajectory probability distributions, an intractable problem for systems with many spatiotemporal degrees of freedom.
In this paper, we propose a method to infer EP in high-dimensional stochastic systems. To do so, we introduce a variational principle that is the nonequilibrium analogue of the well-known Maximum Entropy Principle (MaxEnt) from equilibrium statistical physics. The dual form of this variational principle leads to a tractable, low-dimensional optimization problem. This solution provides an approximation of trajectory-level EP and a lower bound on the average EP, directly from trajectory samples and without explicit knowledge of trajectory probabilities. The dual principle provides a higher-order thermodynamic uncertainty relation (TUR) [14, 15], which we use to derive further simple bounds on EP.
Entropy production.— We consider a nonequilibrium stochastic system measured at discrete time points. The system’s (continuous or discrete) state at any given time is a vector $\bm{x}_{t}=(x_{1,t},x_{2,t},\dots,x_{N,t})$ of dimension $N$. These dimensions may represent physical degrees of freedom, e.g., physical spins or particles in a many-body system. They may also represent population counts (as in a chemical master equation), indicator functions for macrostates, coarse-grained projections, or other types of variables.
The system’s stochastic dynamics over sequences of $T+1$ discrete timesteps are described by a “forward” trajectory distribution $p(\vec{{\bm{x}}})$, where each trajectory $\vec{{\bm{x}}}=({\bm{x}}_{0},{\bm{x}}_{1},\dots,{\bm{x}}_{T})$ is a sequence of $T+1$ states. In general, $T=1$ suffices for stationary (first-order) Markovian systems, while larger $T$ may describe non-Markovian and non-stationary systems. The system is also associated with a “reverse” trajectory distribution $\tilde{p}(\vec{{\bm{x}}})$. The trajectory EP of trajectory $\vec{{\bm{x}}}$ and the average EP across all trajectories are defined as [1]
$\sigma(\vec{{\bm{x}}}):=\ln\frac{p(\vec{{\bm{x}}})}{\tilde{p}(\vec{{\bm{x}}})}\,,\qquad\Sigma:=D(p\|\tilde{p})=\left\langle\ln\frac{p(\vec{{\bm{x}}})}{\tilde{p}(\vec{{\bm{x}}})}\right\rangle_{p}\,,$ (1)
where $\langle\ldots\rangle_{p}=\sum_{\vec{{\bm{x}}}}\ldots p(\vec{{\bm{x}}})$ is the expectation over $p$ and $D(\cdot\|\cdot)$ is the Kullback-Leibler divergence. EP quantifies the violation of time-reversal symmetry. It vanishes in equilibrium when trajectory statistics are indistinguishable under the forward and reverse distributions.
[…]
EP decomposition.— For high-dimensional systems with many degrees of freedom (large $N$) or long memory (large $T$), the set of possible trajectories is exponentially large. In such cases, estimating trajectory probabilities $p(\vec{{\bm{x}}})$ from empirical data becomes statistically infeasible. This prevents estimation of trajectory-level and expected EP by direct evaluation of (1).
We propose a lower bound on EP based on a nonequilibrium generalization of the MaxEnt variational principle. We assume that we have access to $d$ trajectory observables, represented by a vector-valued function $\bm{g}(\vec{{\bm{x}}})\in\mathbb{R}^{d}$. Natural choices of observables include low-order interaction terms, such as 1-point ($x_{i,t}$) and 2-point ($x_{i,t}x_{j,t^{\prime}}$) fluctuating correlations. We do not require antisymmetric observables $\bm{g}(\vec{{\bm{x}}})=-\bm{g}(\reflectbox{$\vec{\reflectbox{${\bm{x}}$}}$})$.
To quantify the EP captured by the temporal irreversibility of observables $\bm{g}$, we minimize the KL divergence to the reverse distribution, with the constraint that the observables $\bm{g}$ have the same expectations as under the forward distribution:
${\Sigma}_{\bm{g}}$ $:=\min_{q}D(q\|\tilde{p})\quad\text{subject to}\quad\left\langle\bm{g}\right\rangle_{q}=\left\langle\bm{g}\right\rangle_{p}.$ (2)
The upper bound $0\leq{\Sigma}_{\bm{g}}\leq{\Sigma}$ follows because the forward distribution $p$ always satisfies the constraints. ${\Sigma}_{\bm{g}}$ vanishes only when $\bm{g}(\vec{{\bm{x}}})$ have the same expectations under forward and backward processes. As we will see, it achieves its maximum ${\Sigma}_{\bm{g}}={\Sigma}$ only when trajectory EP can be expressed as a linear combination of observables $\bm{g}(\vec{{\bm{x}}})$.
Eq. (2) provides a natural and interpretable bound on the EP. However, estimating EP directly from this optimization is not practical in high-dimensional systems, because there is no simple way to evaluate $\tilde{p}(\vec{{\bm{x}}})$, nor to optimize over the exponential (or even infinite) number of parameters that specify the distribution $q$. We solve both problems by exploiting a remarkable information-theoretic duality. As shown in the End Matter, the optimization problem (2) has the dual formulation:
${\Sigma}_{\bm{g}}=\max_{{\bm{\theta}}\in\mathbb{R}^{d}}\left({\bm{\theta}}^{\top}\left\langle\bm{g}\right\rangle_{p}-\ln\langle e^{{\bm{\theta}}^{\top}\bm{g}}\rangle_{\tilde{p}}\right),$ (3)
It may be understood as an optimization over the Lagrangian multipliers ${\bm{\theta}}$ conjugate to constraints on expectations of $\bm{g}$.
Eq. (3) has two important features. First, it avoids estimation of trajectory probabilities, relying only on expectations of $\bm{g}(\vec{{\bm{x}}})$ under $p$ and $e^{{\bm{\theta}}^{\top}\bm{g}(\vec{{\bm{x}}})}$ under $\tilde{p}$, which can be easily computed from empirical samples of observables $\bm{g}(\vec{{\bm{x}}})$.
In addition, Eq. (3) is a low-dimensional optimization problem, involving only as many parameters as observables in $\bm{g}$. For example, for a system with 1000 binary spins measured at two timepoints, Eq. (2) is a constrained optimization with more than $\sim 2^{1000}$ parameters required to specify $q$. For example, it $\bm{g}$ encodes 2-point correlations, then Eq. (3) is an unconstrained convex optimization over $\sim 1000^{2}$ parameters, which is quite feasible for modern computers using standard numerical techniques. In practice, we found that optimizing (3) using standard methods with early stopping on held-out test data led to good performance [16].
So far, we made no assumptions regarding multipartite structure (i.e., only a single variable can change at a time). However, as we show in End Matter, for multipartite dynamics and multipartite observables $\bm{g}(\vec{{\bm{x}}})$, Eq. (3) can be split into $N$ separate optimization problems, each one involving only a proportional fraction of the parameters and observations. This has important numerical advantages, since the computational cost of solving Eq. (3) scales superlinearly with size.
From the perspective of MaxEnt modeling, the distribution $q^{*}$ that optimizes (2) lies on a manifold of probability distributions forming exponential family (see End Matter):
$q_{{\bm{\theta}}}(\vec{{\bm{x}}})=\tilde{p}(\vec{{\bm{x}}})e^{{{\bm{\theta}}}^{\top}\bm{g}(\vec{{\bm{x}}})-\ln\langle e^{{{\bm{\theta}}}^{\top}\bm{g}}\rangle_{\tilde{p}}}.$ (4)
Using this, we can interpret the optimal ${\bm{\theta}}^{*}$ in Eq. (3) as the optimal parameters in the maximum likelihood (ML) problem ${\bm{\theta}}^{*}=\operatorname*{arg\,min}_{{\bm{\theta}}}D(p\|q_{{\bm{\theta}}})$, since $q_{{\bm{\theta}}^{*}}$ satisfies the constraints in (2) and achieves the optimum $q^{*}=q_{{\bm{\theta}}^{*}}$ (see End Matter). This interpretation also leads to the generalized Pythagorean theorem from information geometry [17], which decomposes EP into two nonnegative KL terms:
$\underbrace{D(p\|\tilde{p})}_{{\Sigma}}=$ $\underbrace{D(q^{*}\|\tilde{p})}_{{\Sigma}_{\bm{g}}}+\underbrace{D(p\|q^{*})}_{{\Sigma}^{\perp}_{\bm{g}}}.$ (5)
The term ${\Sigma}^{\perp}_{\bm{g}}:={\Sigma}-{\Sigma}_{\bm{g}}=D(p\|q^{*})$ quantifies the EP that cannot be captured by observables ${\bm{g}}$. It can be expressed as the optimal value of the ML problem, ${\Sigma}^{\perp}_{\bm{g}}=\min_{{\bm{\theta}}\in\mathbb{R}^{d}}D(p\|q_{\bm{\theta}})$. To shed light on this expression, it is useful to define a larger exponential family, $\omega_{f}(\vec{{\bm{x}}}):=\tilde{p}(\vec{{\bm{x}}})e^{f(\vec{{\bm{x}}})}/\langle e^{f}\rangle_{\tilde{p}}$, which contains the forward process as $p=\omega_{\sigma}$. We then write the ML problem as
${\Sigma}^{\perp}_{\bm{g}}=\min_{{\bm{\theta}}\in\mathbb{R}^{d}}D(\omega_{\sigma}\|\omega_{{\bm{\theta}}^{\top}\bm{g}}).$ (6)
It is clear that ${\Sigma}^{\perp}_{\bm{g}}=0$ if and only if $\sigma(\vec{{\bm{x}}})={\bm{\theta}}^{*\top}\bm{g}(\vec{{\bm{x}}})$ for all $\vec{{\bm{x}}}$, meaning that the trajectory EP can be expressed as linear combination of the trajectory observables $\bm{g}$. More generally, the quantity ${\bm{\theta}}^{*\top}\bm{g}(\vec{{\bm{x}}})$ can be considered as the best possible approximation of $\sigma(\vec{{\bm{x}}})$ in the span of $\bm{g}$.
By considering a hierarchy of observables, our approach can define a hierarchical decomposition of EP, similar in spirit to information-geometric decompositions used in equilibrium MaxEnt [17, 18]. For example, if $\bm{g}_{[k]}(\vec{{\bm{x}}})$ specifies the set of $k$-point interactions, then we have the decomposition $0\leq{\Sigma}_{1}\leq{\Sigma}_{2}\leq\dots\leq{\Sigma}$. Here $\Sigma_{m}=D(q^{*}_{m}\|\tilde{p})$ is the contribution from all interactions up to order $m$, defined as the solution of Eq. (2) with expectation constraints on $\langle\bm{g}_{[k]}\rangle$ for $k=1,\dots,m$. Due to the Pythagorean theorem, the EP captured specifically by $k$-point interactions is itself a KL divergence, $\Sigma_{m}-\Sigma_{m-1}=D(q^{*}_{m}\|q^{*}_{m-1})$. In this way, we may derive an interpretable decomposition of how different interactions contribute to EP.
Thermodynamic uncertainty relations (TURs).— Eq. (3) can be interpreted as a TUR that relates EP and fluctuations of trajectory observables [19]. Consider any trajectory observable that can be expressed as a linear combination $o(\vec{{\bm{x}}})={\bm{\theta}}^{\top}\bm{g}(\vec{{\bm{x}}})$ for some ${\bm{\theta}}\in\mathbb{R}^{d}$. The first term in (3) is the expectation of $o(\vec{{\bm{x}}})$, while the second term is the cumulant generating function (CGF) of $o(\vec{{\bm{x}}})$ under $\tilde{p}$. Thus, ${\Sigma}_{\bm{g}}$ bounds the fluctuation-discounted expectation of all such $o(\vec{{\bm{x}}})$, and it can be understood as a higher-order TUR [14] that constrains the variance and the higher-order cumulants. This contrasts with most existing TURs, which quantify fluctuations only using the variance [20, 21, 22].
To relate our approach to variance-based TURs, we expand the CGF as $\ln\langle e^{{\bm{\theta}}^{\top}\bm{g}}\rangle_{\tilde{p}}\approx{\bm{\theta}}^{\top}\left\langle\bm{g}\right\rangle_{\tilde{p}}+{\bm{\theta}}^{\top}\mathsf{K}_{\tilde{p}}{\bm{\theta}}/2$, where $\mathsf{K}_{\tilde{p}}$ is the covariance matrix of $\bm{g}$ under $\tilde{p}$. Plugging into (3) gives the approximate optimizer $\hat{{\bm{\theta}}}$ as the solution to the linear system $\mathsf{K}_{\tilde{p}}\hat{{\bm{\theta}}}=\left\langle\bm{g}\right\rangle_{p}-\left\langle\bm{g}\right\rangle_{\tilde{p}}\equiv\left\langle\bm{g}\right\rangle_{p-\tilde{p}}$. This gives the weaker bound
${\Sigma}_{\bm{g}}\geq\widehat{{\Sigma}}_{\bm{g}}$ $\equiv\left\langle\bm{g}\right\rangle_{p-\tilde{p}}^{\top}\mathsf{K}_{\tilde{p}}^{-1}\left\langle\bm{g}\right\rangle_{p}-\ln\left\langle e^{\left\langle\bm{g}\right\rangle_{p-\tilde{p}}^{\top}\mathsf{K}_{\tilde{p}}^{-1}\bm{g}}\right\rangle_{\tilde{p}}.$ (7)
It also gives the trajectory EP approximation, $\hat{\sigma}_{\bm{g}}(\vec{{\bm{x}}}):=\hat{{\bm{\theta}}}^{\top}\bm{g}(\vec{{\bm{x}}})-\ln\langle e^{{\hat{{\bm{\theta}}}}^{\top}\bm{g}}\rangle_{\tilde{p}}$ which satisfies $\langle\hat{\sigma}_{\bm{g}}\rangle_{p}=\widehat{{\Sigma}}_{\bm{g}}$. The advantage of these estimates is that they do not require optimization, only the solution of a linear system. Moreover, these estimates correspond to a single Newton-Raphson step for maximizing (3) starting from ${\bm{\theta}}=\bm{0}$ (see Supplemental Material (SM) [16]). The bounds become equivalent (${\Sigma}_{\bm{g}}=\widehat{{\Sigma}}_{\bm{g}}$) when $\bm{g}$ has Gaussian statistics.
The optimization-free TUR in (7) may be compared to a quadratic TUR. For stationary systems without odd variables, we find the inequality (see End Matter)
${\Sigma}\geq{\Sigma}_{\bm{g}}^{\textrm{TUR}}:=\frac{1}{2}(\langle\bm{g}\rangle_{p}-\langle\bm{g}\rangle_{\tilde{p}})^{\top}\mathsf{K}_{(p+\tilde{p})/2}^{-1}(\langle\bm{g}\rangle_{p}-\langle\bm{g}\rangle_{\tilde{p}})$ (8)
where $\mathsf{K}_{(p+\tilde{p})/2}$ is the covariance matrix of $\bm{g}$ under the mixture distribution $(p+\tilde{p})/2$. For anti-symmetric observables, Eq. (8) may be considered the discrete-time version of the multidimensional TUR from Ref. [22].
The bounds (7)-(8) may require one to solve linear systems involving very large covariance matrices. For multipartite systems, it is often possible to split these bounds into $N$ bounds that involve smaller matrices (see End Matter).
Related work.— Information-theoretic TURs of the general form (3) were first proposed in [23] and were further explored in [24, 25, 15]. However, these works did not explicitly consider the problem of estimating EP in high-dimensional systems using low-dimensional observables. Several other approaches to EP inference have employed large-deviations theory [26], compression algorithms [27, 28, 29], and deep learning of probability flows [30].
Our Pythagorean theorem (5) is related to previous decompositions of EP in interacting systems [31] and systems with nonconservative forces [32, 15]. Moreover, the general form of Eq. (2) is a Maximum Entropy (MaxEnt) problem over trajectory distributions [33, 34, 35, 36], sometimes called “maximum caliber”  [37]. However, our approach differs from earlier work on MaxEnt. First, we minimize the KL divergence relative to an unknown prior distribution $\tilde{p}$, from which we typically only have some samples. This makes standard techniques and approximation methods for large-scale MaxEnt problems (such as mean-field and Bethe approximations  [38, 39, 6]) not directly applicable. Second, we care not only about the parameters ${\bm{\theta}}$, as is often the case in MaxEnt “inverse problems”, but also about bounding EP by the quantity ${\Sigma}_{\bm{g}}$. Both of these issues are resolved by the dual formulation (3).
Lynn et al. [40, 41, 12] proposed a way to decompose and bound EP using a different information-theoretic optimization. Although originally focused on local EP in multipartite systems, it can be generalized to arbitrary observables $\bm{g}$ as
${\Sigma}_{\bm{g}}^{\textrm{L}}$ $=\min_{q}D(q\|\tilde{q})\;\>\text{subject to}\;\>\langle\bm{g}\rangle_{q}=\langle\bm{g}\rangle_{p},$ (9)
It is usually restricted to stationary system without odd variables, so that $\tilde{q}(\vec{{\bm{x}}})=q(\reflectbox{$\vec{\reflectbox{${\bm{x}}$}}$})$. Like our bound ${\Sigma}_{\bm{g}}$, ${\Sigma}_{\bm{g}}^{\textrm{L}}$ can be used to generate hierarchical decompositions of EP.
Although Eq. (9) resembles Eq. (2), it differs in that it simultaneously optimizes both arguments of the KL divergence, $q$ and $\tilde{q}$. As a result, the resulting optimization problems are dramatically different. In particular, the optimal distribution $q^{*}$ in Eq. (9) is not in an exponential family and it does not have full support in general (see End Matter). Therefore, its dual problem is an optimization over an exponential number of variables, representing the exponential number of Lagrange multipliers required to enforce nonnegativity constraints $q(\vec{{\bm{x}}})\geq 0$ for all $\vec{{\bm{x}}}$. For this reason, at least in their standard form, neither the primal nor dual formulation of Eq. (9) are solvable for large systems, requiring the use of complex dimensionality-reduction techniques [12]. Future work may explore whether Eq. (9) can be solved using other types of optimization techniques.
Examples: Nonequilibrium spin model and Neuropixels dataset.— We illustrate our approach using two examples: a nonequilibrium asymmetric Ising model [42, 43, 6], and in vivo spike data from the Neuropixels Visual Behavior repository [44]. In both systems, we focus on a single time step and binary variables $x_{i,t}\in\{-1,+1\}$, $i\in\{1,\dots,N\}$ with time $t\in\{0,1\}$. We assume that the systems are in steady state, so that samples from $\tilde{p}$ are given by time-reversal of samples from the forward process $p$. For both examples, we optimize Eq. (3) using gradient ascent with Barzilai–Borwein step sizes [45]. We use a validation dataset for early stopping of the optimization, and report EP estimates computed on a held-out test dataset. This avoids overfitting, especially in far-from-equilibrium regimes where many reverse transitions are not sampled. Details of the data, optimization, and analysis are found in the SM [16] and our code repository [46].
The nonequilibrium spin model is specified by transition probabilities
$T(\bm{x}_{1}|\bm{x}_{0})=\frac{1}{N}\sum_{i}\Big{[}W_{i}(\bm{x}_{0})\delta_{\bm{x}_{0},\bm{x}_{1}^{[i]}}+(1-W_{i}(\bm{x}_{0}))\delta_{\bm{x}_{0},\bm{x}_{1}}\Big{]}$
where we introduce the spin-flip operator, $(\bm{x}^{[i]})_{i}=-x_{i}$ and $(\bm{x}^{[i]})_{j}=x_{j}$ for $j\neq i$, and the flip probability for spin $i$,
$W_{i}(\bm{x})$ $=\dfrac{\exp(-\beta x_{i}\sum_{j:j\neq i}w_{ij}x_{j})}{2\cosh(\beta\sum_{j:j\neq i}w_{ij}x_{j})}\,,$ (10)
where $\beta$ is an inverse temperature and $w_{ij}$ are (typically asymmetric) coupling parameters. An advantage of this model is that the steady-state EP has a simple closed-form expression in terms of expectation values, providing the ground truth for our EP estimators. Specifically, ${\Sigma}=\beta\sum_{i\neq j}w_{ij}\langle{(x_{i,1}-x_{i,0})x_{j,0}}\rangle_{\rm st}$, where $\langle\cdot\rangle_{\rm st}$ indicates expectations under the stationary process $p$.
Refer to caption(a)Refer to caption(b)
Figure 1: Disordered nonequilibrium spin model with 1000 spins. (a) Steady-state EP estimation for different inverse temperatures $\beta$. (b) Asymmetry of inferred parameters, shown against the true coupling asymmetries in the model for $\beta=2.5$ ($R^{2}=0.9831$). Estimates are based on $10^{9}$ state transitions sampled by Monte Carlo.
We consider the diluted version of the model [47, 48, 49] with $k$ average neighbors. Here $w_{ij}=c_{ij}z_{ij}/\sqrt{k}$, with binary connections $c_{ij}\sim\textrm{Bernoulli}(k/(N-1))$ and real-valued weights $z_{ij}\sim\mathcal{N}(0,1)$. Our observables of interest are time-lagged correlations,
$g_{ij}(\vec{{\bm{x}}})=(x_{i,1}-x_{i,0})x_{j,0}\qquad\text{for all}\quad i,j\,,$ (11)
with conjugate parameters $\theta_{ij}$. Because the system has multipartite dynamics and observables, Eq. (3) can be decomposed into $N$ independent problems, each involving only those transitions where spin $i$ changes. This dramatically improves the performance scaling of our optimization in system size.
Fig. 1(a) shows the actual and inferred EP at different $\beta$ for $N=1000$, $k=6$. EP increases with $\beta$, and all estimators (${\Sigma}_{\bm{g}}^{\textrm{TUR}}$, $\widehat{{\Sigma}}_{\bm{g}}$, ${\Sigma}_{\bm{g}}$) agree in the near-equilibrium regime of small $\beta$. Moreover, ${\Sigma}_{\bm{g}}$ provides a tight bound on EP even in the far-from-equilibrium regime of large $\beta$. The gap between ${\Sigma}_{\bm{g}}$ and $\widehat{{\Sigma}}_{\bm{g}}$ indicates onset of highly non-Gaussian statistics from $\beta\approx 2$.
As discussed near Eq. (6), ${\Sigma}_{\bm{g}}\approx{\Sigma}$ implies that the trajectory EP $\sigma(\vec{{\bm{x}}})$ can be closely approximated by a linear combination of observables $g_{ij}$. This is a nontrivial result even for this model, since $\sigma(\vec{{\bm{x}}})$ contains contributions from the difference of steady-state probabilities, which is not known in closed form. As we show in the SM, our optimal parameters ${\bm{\theta}}^{*}$ can be used to infer the asymmetry of the coupling constants as $\theta_{ij}^{*}-\theta_{ji}^{*}\approx\beta(w_{ij}-w_{ji})$. This relation is verified in Fig. 1(b).
In the SM, we also consider a different family of Gaussian coupling matrices, for which this model exhibits a nonequilibrium phase transition [6]. We show that our approach accurately infers EP in this regime, which has been shown to be difficult for classical approximations [50].
(a)Refer to captionRefer to caption(b)
Figure 2: (a) Estimated EP per expected number of spikes $R$, in the Neuropixels Visual Behavior dataset for three conditions. Error bars indicate standard error of the mean. (b) Example of the values of inferred coupling parameters $\theta^{*}_{ij}$ from one trial in the active condition selecting the 200 most active neurons, grouped by visual brain area. To improve visualization, lower triangle shows $\theta_{ji}^{*}\equiv-\theta_{ij}^{*}$ for $i<j$.
Next, as an example application to biological data, we estimate EP in the Visual Behavior Neuropixels dataset [44], containing spike-train recordings from 81 mice across 103 sessions. The dataset includes spiking activity from multiple brain regions, including visual cortical areas (VISp, VISl, VISal, VISrl, VISam and VISpm) and subcortical structures. We analyze data from visual areas during three conditions: (1) active behavior (visual change detection task), (2) passive replay (same stimuli but without task engagement), and (3) Gabor (receptive field characterization with Gabor stimuli and full-field flashes). We discretize spike trains into temporal bins of length $\Delta t=10$ ms and verify that almost all bins contain no more than one spike. In this dataset, neurons can update in a parallel (non-multipartite) manner.
Our observables are defined as
$g_{ij}(\vec{{\bm{x}}})=x_{i,1}\,x_{j,0}-x_{i,0}\,x_{j,1}\qquad\text{for all}\quad i<j\,,$ (12)
which naturally represent time-lagged correlations in non-multipartite dynamics. They capture the antisymmetric part of  (11) and give more robust numerical results on this dataset.
We calculate EP values across different conditions and different system sizes, $N$. For each combination of size, condition, and recording session, we randomly selected 10 subsets of neurons and estimated ${\Sigma}_{\bm{g}}$ for each subset. For improved comparison, we normalize EP in each condition by the expected number of spikes per bin, $R=\sum_{i}\left(1+\langle x_{i,0}\rangle_{\rm st}\right)/2$. EP is here interpreted as a statistical measure of time-irreversibility, not energetic dissipation.
Fig. 2(a) shows that EP grows superlinearly with size (note that $R\propto N$) and that the active condition is associated with the largest normalized EP. Fig. 2(b) shows the matrix of inferred parameters $\theta_{ij}^{*}$, grouped by visual brain area. This shows that our method infers the pairwise coupling coefficients that determine temporal irreversibility in a neural system, revealing a clustered organization that is meaningfully aligned with anatomical brain regions.
Author Contributions
A.K. and M.A. contributed equally to this work.
Acknowledgements
Acknowledgements.
M.A. is partly supported by John Templeton Foundation (grant 62828) and Grant PID2023-146869NA-I00 funded by MICIU/AEI/10.13039/501100011033 and cofunded by the European Union, and supported by the Basque Government through the BERC 2022-2025 program and by the Spanish State Research Agency through BCAM Severo Ochoa excellence accreditation CEX2021-01142-S funded by MICIU/AEI/10.13039/501100011033. S. I. is supported by JSPS KAKENHI Grants No. 22H01141, No. 23H00467, and No. 24H00834, JST ERATO Grant No. JPMJER2302, and UTEC-UTokyo FSI Research Grant Program. A.K. is partly supported by John Templeton Foundation (grant 62828) and by the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie Grant Agreement No. 101068029.
I Data Availability Statement
The code used in this study is available at a GitHub repository [46].