### v2

# Inferring entropy production in many-body systems using nonequilibrium MaxEnt

Abstract
We propose a method for inferring entropy production (EP) in high-dimensional stochastic systems, including many-body systems and non-Markovian systems with long memory. Standard techniques for estimating EP become intractable in such systems due to computational and statistical limitations. We infer trajectory-level EP and lower bounds on average EP by exploiting a nonequilibrium analogue of the Maximum Entropy principle, along with convex duality. Our approach uses only samples of trajectory observables (such as spatiotemporal correlation functions). It does not require reconstruction of high-dimensional probability distributions or rate matrices, nor any special assumptions such as discrete states or multipartite dynamics. It may be used to compute a hierarchical decomposition of EP, reflecting contributions from different kinds of interactions, and it has an intuitive physical interpretation as a thermodynamic uncertainty relation. We demonstrate its numerical performance on a disordered nonequilibrium spin model with 1000 spins and a large neural spike-train dataset.
The central quantity of interest in nonequilibrium thermodynamics is entropy production (EP) [1]. EP characterizes departure from thermodynamic equilibrium in microscopic physical systems, where it quantifies the dissipation of thermodynamic free energy. More generally, EP provides an information-theoretic measure of temporal irreversibility, including in meso- and macroscopic systems [2, 3, 4, 5]. There is great interest in quantifying EP in high-dimensional complex systems with many degrees of freedom and/or long non-Markovian memory, such as nonequilibrium disordered networks [6], biological active matter [7, 8], and neural systems [9, 10, 11, 12, 13], but this is computationally and statistically difficult. Standard methods require reconstruction of trajectory probability distributions, an intractable problem for systems with many spatiotemporal degrees of freedom.
In this paper, we propose a method to infer EP in high-dimensional stochastic systems. To do so, we introduce a variational principle that is the nonequilibrium analogue of the well-known Maximum Entropy Principle (MaxEnt) from equilibrium statistical physics. The dual form of this variational principle leads to a tractable optimization problem whose solution provides an approximation of trajectory-level EP and a lower bound on the average EP — directly from trajectory samples and without explicit knowledge of trajectory probabilities. The dual principle also provides a higher-order thermodynamic uncertainty relation (TUR) [14, 15], which we use to derive further simple bounds on EP.
Entropy production.— We consider a nonequilibrium stochastic system measured at discrete time points. The system’s (continuous or discrete) state at any given time is a vector $\bm{x}_{t}=(x_{1,t},x_{2,t},\dots,x_{N,t})$ of dimension $N$. These dimensions may represent physical degrees of freedom, e.g., physical spins or particles in a many-body system. They may also represent population counts (as in a chemical master equation), indicator functions for macrostates, coarse-grained projections, or other types of variables.
The system’s stochastic dynamics over sequences of $T+1$ discrete timesteps are described by a “forward” trajectory distribution $p(\vec{{\bm{x}}})$, where each trajectory $\vec{{\bm{x}}}=({\bm{x}}_{0},{\bm{x}}_{1},\dots,{\bm{x}}_{T})$ is a sequence of $T+1$ states. In general, $T=1$ suffices for stationary first-order-Markovian systems, while larger $T$ may describe non-Markovian and non-stationary systems. The system is also associated with a “reverse” trajectory distribution $\tilde{p}(\vec{{\bm{x}}})$. The trajectory EP of trajectory $\vec{{\bm{x}}}$ and the average EP across all trajectories are defined as [1]
$\sigma(\vec{{\bm{x}}}):=\ln\frac{p(\vec{{\bm{x}}})}{\tilde{p}(\vec{{\bm{x}}})}\,,\qquad\Sigma:=D(p\|\tilde{p})=\left\langle\ln\frac{p(\vec{{\bm{x}}})}{\tilde{p}(\vec{{\bm{x}}})}\right\rangle_{p}\,,$ (1)
where $\langle\ldots\rangle_{p}=\sum_{\vec{{\bm{x}}}}\ldots p(\vec{{\bm{x}}})$ is the expectation over $p$ and $D(\cdot\|\cdot)$ is the Kullback-Leibler (KL) divergence. EP quantifies the violation of time-reversal symmetry, and it vanishes in equilibrium when trajectory statistics are indistinguishable under the forward and reverse distributions.
[…]
EP decomposition.— For high-dimensional systems with many degrees of freedom (large $N$) or long memory (large $T$), the set of possible trajectories is exponentially large. In such cases, estimating trajectory probabilities $p(\vec{{\bm{x}}})$ from empirical data becomes statistically infeasible. This prevents estimation of trajectory-level and expected EP by direct evaluation of (1).
We propose a lower bound on EP based on a nonequilibrium generalization of the MaxEnt variational principle. We assume that we have access to $d$ trajectory observables, represented by a vector-valued function $\bm{g}(\vec{{\bm{x}}})\in\mathbb{R}^{d}$. Natural choices of observables include low-order interaction terms, such as 1-point ($x_{i,t}$) and 2-point ($x_{i,t}x_{j,t^{\prime}}$) correlations. We do not require antisymmetric observables $\bm{g}(\vec{{\bm{x}}})=-\bm{g}(\reflectbox{$\vec{\reflectbox{${\bm{x}}$}}$})$.
To quantify the EP captured by the temporal irreversibility of observables $\bm{g}$, we find the distribution that has the smallest KL divergence to the reverse process, subject to having the same observable expectations as the forward process:
${\Sigma}_{\bm{g}}$ $:=\min_{q}D(q\|\tilde{p})\quad\text{subject to}\quad\left\langle\bm{g}\right\rangle_{q}=\left\langle\bm{g}\right\rangle_{p}.$ (2)
${\Sigma}_{\bm{g}}$ is always nonnegative, and it vanishes only when the expectation of $\bm{g}$ is the same under the forward and reverse processes. Furthermore, it is straightforward to show that ${\Sigma}_{\bm{g}}\leq D(p_{G}\|\tilde{p}_{G})\leq{\Sigma}$, where $p_{G}(\bm{g}^{\prime})=\sum_{\vec{{\bm{x}}}}p(\vec{{\bm{x}}})\delta_{\bm{g}^{\prime},\bm{g}(\vec{{\bm{x}}})}$ and $\tilde{p}_{G}(\bm{g}^{\prime})=\sum_{\vec{{\bm{x}}}}\tilde{p}(\vec{{\bm{x}}})\delta_{\bm{g}^{\prime},\bm{g}(\vec{{\bm{x}}})}$ indicate the distributions of $\bm{g}$ under the forward and backward processes (see Supplemental Material (SM) [16]). In general, estimating $D(p_{G}\|\tilde{p}_{G})$ is as difficult as estimating ${\Sigma}$, since it requires inference of the high-dimensional distributions $p_{G}$ and $\tilde{p}_{G}$.
On the other hand, by exploiting a remarkable information-theoretic duality, we may calculate ${\Sigma}_{\bm{g}}$ without explicitly inferring any probability distributions. Specifically, as shown in the End Matter, the optimization problem (2) has the dual formulation
${\Sigma}_{\bm{g}}=\max_{{\bm{\theta}}\in\mathbb{R}^{d}}\left({\bm{\theta}}^{\top}\left\langle\bm{g}\right\rangle_{p}-\ln\langle e^{{\bm{\theta}}^{\top}\bm{g}}\rangle_{\tilde{p}}\right).$ (3)
Eq. (3) is a tractable convex optimization problem over the Lagrangian multipliers ${\bm{\theta}}$ that enforce the $d$ expectation constraints on $\bm{g}$. Importantly, the objective depends only on the expectations of $\bm{g}(\vec{{\bm{x}}})$ under $p$ and $e^{{\bm{\theta}}^{\top}\bm{g}(\vec{{\bm{x}}})}$ under $\tilde{p}$, which can be easily computed from empirical samples of the observables.
As an example, consider a system with 1000 binary spins measured at two timepoints. A naive estimate of ${\Sigma}$ would require inference of $\sim 2^{1000}$ probabilities. Instead, with $\bm{g}$ encoding 2-point correlations, Eq. (3) results in a tractable unconstrained optimization problem over $\sim 1000^{2}$ variables.
So far, we made no assumptions regarding multipartite structure (i.e., only a single variable can change at a time). However, as we show in End Matter, for multipartite dynamics and multipartite observables, Eq. (3) can be split into $N$ separate optimization problems, each one involving only a proportional fraction of the variables and observations. This can dramatically reduce the amount of computation and memory required to numerically solve Eq. (3).
The distribution $q^{*}$ that optimizes (2) lies in the exponential family,
$q_{{\bm{\theta}}}(\vec{{\bm{x}}})=\tilde{p}(\vec{{\bm{x}}})e^{{{\bm{\theta}}}^{\top}\bm{g}(\vec{{\bm{x}}})-\ln\langle e^{{{\bm{\theta}}}^{\top}\bm{g}}\rangle_{\tilde{p}}}.$ (4)
We can interpret the optimal ${\bm{\theta}}^{*}$ in Eq. (3) as the optimal parameters in the maximum likelihood (ML) problem ${\bm{\theta}}^{*}=\operatorname*{arg\,min}_{{\bm{\theta}}}D(p\|q_{{\bm{\theta}}})$, since $q_{{\bm{\theta}}^{*}}=q^{*}$ satisfies the constraints in (2) and achieves the optimum objective value (see End Matter). This interpretation also leads to the information-geometric Pythagorean theorem [17], which decomposes EP into two nonnegative KL terms:
$\underbrace{D(p\|\tilde{p})}_{{\Sigma}}=$ $\underbrace{D(q^{*}\|\tilde{p})}_{{\Sigma}_{\bm{g}}}+\underbrace{D(p\|q^{*})}_{{\Sigma}^{\perp}_{\bm{g}}}.$ (5)
The term ${\Sigma}^{\perp}_{\bm{g}}:={\Sigma}-{\Sigma}_{\bm{g}}=D(p\|q^{*})$ quantifies the EP not captured by ${\Sigma}_{\bm{g}}$. It can be expressed in terms of a ML problem, ${\Sigma}^{\perp}_{\bm{g}}=\min_{{\bm{\theta}}\in\mathbb{R}^{d}}D(p\|q_{\bm{\theta}})$. To interpret this expression, we define a larger exponential family $\omega_{f}(\vec{{\bm{x}}}):=\tilde{p}(\vec{{\bm{x}}})e^{f(\vec{{\bm{x}}})}/\langle e^{f}\rangle_{\tilde{p}}$ that contains the forward trajectory distribution $p(\vec{{\bm{x}}})=\omega_{\sigma}(\vec{{\bm{x}}})$. The ML problem can be written as
${\Sigma}^{\perp}_{\bm{g}}=\min_{{\bm{\theta}}\in\mathbb{R}^{d}}D(\omega_{\sigma}\|\omega_{{\bm{\theta}}^{\top}\bm{g}}).$ (6)
From this expression, it is clear that ${\Sigma}^{\perp}_{\bm{g}}=0$ (and ${\Sigma}_{\bm{g}}={\Sigma}$) if and only if trajectory EP is a linear combination of the trajectory observables, i.e., $\sigma(\vec{{\bm{x}}})={\bm{\theta}}^{*\top}\bm{g}(\vec{{\bm{x}}})$ for all $\vec{{\bm{x}}}$. When ${\Sigma}^{\perp}_{\bm{g}}>0$, ${\bm{\theta}}^{*\top}\bm{g}(\vec{{\bm{x}}})$ provides the best possible approximation of $\sigma(\vec{{\bm{x}}})$ in the span of $\bm{g}$.
Eq. (4) implies that the conditional distribution of the optimal $q^{*}$ obeys $q^{*}_{\vec{X}|G}(\vec{{\bm{x}}}|\bm{g}^{\prime})=\tilde{p}_{\vec{X}|G}(\vec{{\bm{x}}}|\bm{g}^{\prime})$, with conditional probabilities given by Bayes’ rule, $p_{\vec{X}|G}(\vec{{\bm{x}}}|\bm{g}^{\prime})=p(\vec{{\bm{x}}})\delta_{\bm{g}^{\prime},\bm{g}(\vec{{\bm{x}}})}/p_{G}(\bm{g}^{\prime})$. Using the chain rule of KL divergence, we may further decompose ${\Sigma}^{\perp}_{\bm{g}}$ into
${\Sigma}^{\perp}_{\bm{g}}=D(p_{G}\|q^{*}_{G})+D(p_{\vec{X}|G}\|\tilde{p}_{\vec{X}|G})\,.$ (7)
The first term in Eq. (7) vanishes when $q^{*}_{G}=p_{G}$; it is the additional EP that could be inferred by constraining all higher-order statistics (not just the mean) in Eq. (2). The second term is the EP that cannot be inferred from any statistics of $\bm{g}$.
Finally, our method supports a hierarchical decomposition of EP, where contributions from interactions of increasing order (e.g., pairwise, triplets) are successively added, resulting in a structure $0\leq{\Sigma}_{1}\leq{\Sigma}_{2}\leq\dots\leq{\Sigma}$, with $\Sigma_{k}=D(q^{*}_{k}\|\tilde{p})$ capturing contributions from functions $\bm{g}(\vec{{\bm{x}}})$ containing interactions up to order $k$. This structure mirrors information-geometric decompositions in equilibrium MaxEnt [17, 18] and allows EP to be broken down into interpretable components, each satisfying $\Sigma_{k}-\Sigma_{k-1}=D(q^{*}_{k}\|q^{*}_{k-1})$.
Thermodynamic uncertainty relations (TURs).— Eq. (3) can be interpreted as a TUR that relates EP and fluctuations of trajectory observables [19]. Consider any trajectory observable that can be expressed as a linear combination $o(\vec{{\bm{x}}})={\bm{\theta}}^{\top}\bm{g}(\vec{{\bm{x}}})$ for some ${\bm{\theta}}\in\mathbb{R}^{d}$. The first term in (3) is the expectation of $o(\vec{{\bm{x}}})$, while the second term is the cumulant generating function (CGF) of $o(\vec{{\bm{x}}})$ under $\tilde{p}$. Thus, ${\Sigma}_{\bm{g}}$ bounds the “fluctuation discounted” expectation of all such $o(\vec{{\bm{x}}})$, giving a higher-order TUR [14] that constrains all cumulants of trajectory observables. This contrasts with quadratic TURs, which constrain only the mean and variance [20, 21, 22].
To relate our approach to quadratic TURs, we expand the CGF as $\ln\langle e^{{\bm{\theta}}^{\top}\bm{g}}\rangle_{\tilde{p}}\approx{\bm{\theta}}^{\top}\left\langle\bm{g}\right\rangle_{\tilde{p}}+{\bm{\theta}}^{\top}\mathsf{K}_{\tilde{p}}{\bm{\theta}}/2$, where $\mathsf{K}_{\tilde{p}}$ is the covariance matrix of $\bm{g}$ under $\tilde{p}$. Plugging into (3) gives the approximate optimizer $\hat{{\bm{\theta}}}$ as the solution to the linear system $\mathsf{K}_{\tilde{p}}\hat{{\bm{\theta}}}=\left\langle\bm{g}\right\rangle_{p}-\left\langle\bm{g}\right\rangle_{\tilde{p}}\equiv\left\langle\bm{g}\right\rangle_{p-\tilde{p}}$. This gives the weaker bound
${\Sigma}_{\bm{g}}\geq\widehat{{\Sigma}}_{\bm{g}}$ $:=\left\langle\bm{g}\right\rangle_{p-\tilde{p}}^{\top}\mathsf{K}_{\tilde{p}}^{-1}\left\langle\bm{g}\right\rangle_{p}-\ln\left\langle e^{\left\langle\bm{g}\right\rangle_{p-\tilde{p}}^{\top}\mathsf{K}_{\tilde{p}}^{-1}\bm{g}}\right\rangle_{\tilde{p}}.$ (8)
$\widehat{{\Sigma}}_{\bm{g}}$ can be shown to correspond to a single Newton-Raphson step for maximizing (3) starting from ${\bm{\theta}}=\bm{0}$ [16]; it has the advantage of not requiring optimization, only the solution a linear system. The bound ${\Sigma}_{\bm{g}}\geq\widehat{{\Sigma}}_{\bm{g}}$ becomes tight when $\bm{g}$ has Gaussian statistics.
We may also compare our results to a quadratic TUR. For stationary systems without odd variables, we find the inequality (see End Matter)
${\Sigma}\geq{\Sigma}_{\bm{g}}^{\textrm{TUR}}:=\frac{1}{2}(\langle\bm{g}\rangle_{p}-\langle\bm{g}\rangle_{\tilde{p}})^{\top}\mathsf{K}_{(p+\tilde{p})/2}^{-1}(\langle\bm{g}\rangle_{p}-\langle\bm{g}\rangle_{\tilde{p}})$ (9)
where $\mathsf{K}_{(p+\tilde{p})/2}$ is the covariance matrix of $\bm{g}$ under the mixture distribution $(p+\tilde{p})/2$. For antisymmetric observables, Eq. (9) may be considered the discrete-time version of the multidimensional TUR from Ref. [22].
The bounds (8)-(9) may require solving linear systems involving very large covariance matrices. For multipartite systems, it is often possible to split these bounds into $N$ bounds that involve smaller matrices (see End Matter).
Related work.— Eq. (3) is closely related to the Donsker-Varadhan representation from large deviations theory [23], previously used to estimate information-theoretic quantities in machine learning [24] and thermodynamic inference [25, 26, 27]. However, these works did not consider the problem of thermodynamic inference in high-dimensional systems using multivariate observables. Other approaches to EP inference have employed compression algorithms [28, 29, 30] and deep learning of probability flows [31].
Our Pythagorean theorem (5) is related to previous decompositions of EP in interacting systems [32] and systems with nonconservative forces [33, 34, 15]. Moreover, the general form of Eq. (2) is a Maximum Entropy (MaxEnt) problem over trajectory distributions [35, 36, 37, 38], sometimes called “maximum caliber” (MaxCal) [39]. Related techniques have been used to infer models from which the entropy flow can be estimated, though without ensuring a lower bound [40].
Our approach differs in several other ways from earlier work on MaxEnt and MaxCal. First, we minimize the KL divergence relative to an unknown prior distribution $\tilde{p}$, from which we typically only have some samples. This makes standard approximation methods for large-scale MaxEnt problems (such as mean-field and Bethe approximations [41, 42, 6]) not directly applicable. Second, we care not only about the parameters ${\bm{\theta}}$, as is typical of MaxEnt “inverse problems”, but also about bounding EP by the quantity ${\Sigma}_{\bm{g}}$. Both of these issues are resolved by the dual formulation (3).
Lynn et al. [43, 44, 12] proposed a way to decompose and bound EP using a different information-theoretic optimization. Although originally focused on local EP in multipartite stationarity systems, it can be generalized to arbitrary observables $\bm{g}$ and non-multipartite systems as
${\Sigma}_{\bm{g}}^{\textrm{L}}$ $=\min_{q,\tilde{q}}D(q\|\tilde{q})\;\>\text{subject to}\;\>\langle\bm{g}\rangle_{q}=\langle\bm{g}\rangle_{p}.$ (10)
Here we use assumption of stationarity and no odd variables, so that $\tilde{q}(\vec{{\bm{x}}})=q(\reflectbox{$\vec{\reflectbox{${\bm{x}}$}}$})$. Like our bound ${\Sigma}_{\bm{g}}$, ${\Sigma}_{\bm{g}}^{\textrm{L}}$ can be used to generate hierarchical decompositions of EP.
Although Eq. (10) resembles Eq. (2), it differs in that it simultaneously optimizes both arguments of the KL divergence, $q$ and $\tilde{q}$. As we discuss in the End Matter, the optimal distribution $q^{*}$ in Eq. (10) is not in an exponential family and in general it lacks full support. For this reason, Eq. (10) does not have a tractable dual expression (i.e., analogous to our dual (3)), and there is no straightforward way to scale ${\Sigma}_{\bm{g}}^{\textrm{L}}$ to large systems.
Examples: Nonequilibrium spin model and Neuropixels dataset.— We illustrate our approach using two examples: a nonequilibrium kinetic Ising model [45, 46, 6], and in vivo spike data from the Neuropixels Visual Behavior repository [47]. In both systems, we focus on a single time step and binary variables $x_{i,t}\in\{-1,+1\}$, $i\in\{1,\dots,N\}$ with time $t\in\{0,1\}$, and assume stationarity.
To calculate ${\Sigma}_{\bm{g}}$, we optimize Eq. (3) using gradient ascent with Barzilai–Borwein step sizes [48]. We perform early-stopping using held-out validation data, which avoids overfitting even in the far-from-equilibrium regime where many reverse transitions are not sampled. Reported EP estimates are computed on held-out test data. Details of the data generation, optimization, and analysis are found in the SM [16] and our code repository [49].
The nonequilibrium spin model is specified by transition probabilities
$T(\bm{x}_{1}|\bm{x}_{0})=\frac{1}{N}\sum_{i}\Big{[}W_{i}(\bm{x}_{0})\delta_{\bm{x}_{0},\bm{x}_{1}^{[i]}}+(1-W_{i}(\bm{x}_{0}))\delta_{\bm{x}_{0},\bm{x}_{1}}\Big{]}$
Here we introduce the spin-flip operator, $(\bm{x}^{[i]})_{i}=-x_{i}$ and $(\bm{x}^{[i]})_{j}=x_{j}$ for $j\neq i$. The flip probability for spin $i$ is
$W_{i}(\bm{x})$ $=\dfrac{\exp(-\beta x_{i}\sum_{j:j\neq i}w_{ij}x_{j})}{2\cosh(\beta\sum_{j:j\neq i}w_{ij}x_{j})}\,,$ (11)
where $\beta$ is an inverse temperature and $w_{ij}$ are (typically asymmetric) coupling parameters. We consider the diluted version of the model [50, 51, 52] with $k$ average neighbors. Here $w_{ij}=c_{ij}z_{ij}/\sqrt{k}$, with binary connections $c_{ij}\sim\textrm{Bernoulli}(k/(N-1))$ and real-valued weights $z_{ij}\sim\mathcal{N}(0,1)$.
The steady-state EP in this model has a simple closed-form expression, ${\Sigma}=\beta\sum_{i\neq j}w_{ij}\langle{(x_{i,1}-x_{i,0})x_{j,0}}\rangle_{\rm st}$, where $\langle\cdot\rangle_{\rm st}$ indicates expectations under the stationary process $p$. This provides ground truth to evaluate our estimators.
Refer to caption(a)Refer to caption(b)
Figure 1: Disordered nonequilibrium spin model with 1000 spins. (a) Steady-state EP estimates for different inverse temperatures $\beta$. (b) Asymmetry of inferred parameters, shown against the true coupling asymmetries in the model for $\beta=2.5$ ($R^{2}=0.9831$). Estimates are based on $10^{9}$ state transitions sampled by Monte Carlo.
Our observables of interest are time-lagged correlations,
$g_{ij}(\vec{{\bm{x}}})=(x_{i,1}-x_{i,0})x_{j,0}\qquad\text{for all}\quad i,j\,,$ (12)
with conjugate parameters $\theta_{ij}$. Because the system has multipartite dynamics and observables, Eq. (3) can be decomposed into $N$ independent problems, each involving only those transitions where spin $i$ changes, improving computational performance for large systems.
Fig. 1(a) shows the actual and inferred EP at different $\beta$ for $N=1000$, $k=6$. EP increases with $\beta$, and all estimators (${\Sigma}_{\bm{g}}^{\textrm{TUR}}$, $\widehat{{\Sigma}}_{\bm{g}}$, ${\Sigma}_{\bm{g}}$) agree in the near-equilibrium regime of small $\beta$. Importantly, ${\Sigma}_{\bm{g}}$ provides a tight bound on EP even in the far-from-equilibrium regime of large $\beta$. The gap between ${\Sigma}_{\bm{g}}$ and $\widehat{{\Sigma}}_{\bm{g}}$ indicates the onset of highly non-Gaussian statistics from $\beta\approx 2$.
As discussed near Eq. (6), ${\Sigma}_{\bm{g}}\approx{\Sigma}$ implies that the trajectory EP $\sigma(\vec{{\bm{x}}})$ can be closely approximated by a linear combination of observables $g_{ij}$. As shown in the SM, our optimal parameters ${\bm{\theta}}^{*}$ can be used to infer the asymmetry of the coupling constants, $\theta_{ij}^{*}-\theta_{ji}^{*}\approx\beta(w_{ij}-w_{ji})$. This relation is verified in Fig. 1(b).
In the SM, we also consider a different family of coupling matrices, for which this model exhibits a nonequilibrium phase transition [6]. We show that our approach accurately infers EP in this regime, a difficult task for existing approximations [53].
(a)Refer to captionRefer to caption(b)
Figure 2: (a) Estimated EP per expected number of spikes $R$, in the Neuropixels Visual Behavior dataset for three conditions. Error bars indicate standard error of the mean. (b) A sample of inferred coupling coefficients $\theta^{*}_{ij}$ grouped by visual area. Here we select 200 neurons with highest firing rate from an active trial. To improve visualization, lower triangle shows $\theta_{ji}^{*}\equiv-\theta_{ij}^{*}$ for $i<j$.
Next, as an example application to biological data, we estimate EP in the Visual Behavior Neuropixels dataset [47], containing spike-train recordings from 81 mice across 103 sessions. EP is here interpreted as a statistical measure of time-irreversibility, not energetic dissipation [13]. The dataset includes spiking activity from multiple brain regions, including visual cortical areas (VISp, VISl, VISal, VISrl, VISam and VISpm) and subcortical structures. We analyze data from visual areas during three conditions: active behavior (visual change detection task), passive replay (same stimuli but without task engagement), and Gabor (receptive field characterization with Gabor stimuli and full-field flashes). We discretize spike trains into temporal bins of length $\Delta t=10$ ms and verify that most bins contain no more than one spike. We note that neurons can update in a parallel (non-multipartite) manner.
Our observables are defined as
$g_{ij}(\vec{{\bm{x}}})=x_{i,1}\,x_{j,0}-x_{i,0}\,x_{j,1}\qquad\text{for all}\quad i<j\,,$ (13)
which naturally represent time-lagged correlations in non-multipartite dynamics. They capture the antisymmetric part of (12) and give more robust numerical results on this dataset.
We calculate EP estimates across different conditions and system sizes $N$. For each combination of size, condition, and recording session, we randomly select 10 subsets of neurons and estimate ${\Sigma}_{\bm{g}}$ for each subset. For improved comparison, we normalize EP in each condition by the expected number of spikes per bin, $R=\sum_{i}\left(1+\langle x_{i,0}\rangle_{\rm st}\right)/2$.
Fig. 2(a) shows that EP grows superlinearly with size (note that $R\propto N$). It also shows that the active condition is associated with the largest normalized EP. Fig. 2(b) illustrates the matrix of inferred parameters $\theta_{ij}^{*}$ for 200 neurons from an active trial, grouped by visual brain area. These parameters can be understood as the coupling coefficients that best describe temporal irreversibility of the in vivo neural dynamics. Notably, they reveal a clustered organization that is aligned with anatomical brain regions.
Author Contributions
A.K. and M.A. contributed equally to this work.
Acknowledgements
Acknowledgements.
M.A. is partly supported by John Templeton Foundation (grant 62828) and Grant PID2023-146869NA-I00 funded by MICIU/AEI/10.13039/501100011033 and cofunded by the European Union, and supported by the Basque Government through the BERC 2022-2025 program and by the Spanish State Research Agency through BCAM Severo Ochoa excellence accreditation CEX2021-01142-S funded by MICIU/AEI/10.13039/501100011033. S. I. is supported by JSPS KAKENHI Grants No. 22H01141, No. 23H00467, and No. 24H00834, JST ERATO Grant No. JPMJER2302, and UTEC-UTokyo FSI Research Grant Program. A.K. is partly supported by John Templeton Foundation (grant 62828) and by the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie Grant Agreement No. 101068029.
I Data Availability Statement
The code used in this study is available at a GitHub repository [49].