### v3

# Inferring entropy production in many-body systems using nonequilibrium MaxEnt

Abstract
We propose a method for inferring entropy production (EP) in high-dimensional stochastic systems, including many-body systems and non-Markovian systems with long memory. Standard techniques for estimating EP become intractable in such systems due to computational and statistical limitations. We infer trajectory-level EP and lower bounds on average EP by exploiting a nonequilibrium analogue of the Maximum Entropy principle, along with convex duality. Our approach uses only samples of trajectory observables, such as spatiotemporal correlations. It does not require reconstruction of high-dimensional probability distributions or rate matrices, nor impose any special assumptions such as discrete states or multipartite dynamics. In addition, it may be used to compute a hierarchical decomposition of EP, reflecting contributions from different interaction orders, and it has an intuitive physical interpretation as a “thermodynamic uncertainty relation.” We demonstrate its numerical performance on a disordered nonequilibrium spin model with 1000 spins and a large neural spike-train dataset.
The central quantity of interest in nonequilibrium thermodynamics is entropy production (EP) [1]. In microscopic physical systems, EP characterizes departure from thermodynamic equilibrium and quantifies the dissipation of thermodynamic free energy. More generally, EP provides an information-theoretic measure of temporal irreversibility, including in meso- and macroscopic systems [2, 3, 4, 5].
Recently, there has been growing interest in thermodynamic inference [6], that is, the problem of estimating EP from empirical measurements of a stochastic system. Various methods have been developed for inferring EP from partial [7, 8, 9, 10, 11, 12, 13, 14] information. For instance, the celebrated thermodynamic uncertainty relation (TUR) [15, 16, 17, 18] bounds EP in terms of the mean and variance of a single current. Other techniques relate EP to the statistics of waiting times [19, 20, 21, 22], observed transitions [23, 24, 11, 12], and counting observables [25]. Many of these methods are designed to infer underlying dissipation from a small number of coarse-grained observables.
In this Letter, we consider thermodynamic inference in the high-dimensional setting, where a large number of observables are available. This is particularly relevant for multivariate measurements of complex systems — such as nonequilibrium disordered networks [26], biological active matter [27, 28], and neural systems [29, 30, 31, 32, 33] — that have many degrees of freedom and/or long non-Markovian memory. A naïve approach requires the estimation of trajectory probability distributions, but in high dimensions, this is usually statistically and numerically infeasible.
To address this challenge, we propose to infer EP using an information-theoretic variational principle, which can be understood as the nonequilibrium analogue of the Maximum Entropy Principle (MaxEnt) in statistical physics. Our variational principle has a simple physical interpretation: it quantifies the irreversibility captured by the expectations of some (possibly large) number of trajectory observables. Importantly, this variational principle has a dual form that leads to a tractable convex optimization problem. This optimization problem gives a lower bound on EP and an estimate of the fluctuating trajectory-level EP — directly from trajectory samples and without explicit use of trajectory probabilities.
As we show, our bound can also be understood as a higher-order thermodynamic uncertainty relation (TUR) [34, 35], which we use to derive further simple bounds on EP. Finally, we show that in multipartite systems, our variational principle can be split into smaller subproblems, greatly improving the performance scaling of our method with system size.
[…]
The trajectory EP of trajectory $\bm{x}$}} and the average EP across all trajectories are defined in terms of symmetry breaking between forward and reverse distributions [1],
$\sigma({{\bm{x}}})\vcentcolon=\ln\frac{p({{\bm{x}}})}{\tilde{p}({{\bm{x}}})}\,,\qquad\Sigma\vcentcolon=D(p\|\tilde{p})=\left\langle\ln\frac{p({{\bm{x}}})}{\tilde{p}({{\bm{x}}})}\right\rangle_{p}\,$, (1)
where $\langle\ldots\rangle_{p}=\sum_{{{\bm{x}}}}\ldots p({{\bm{x}}})$ is the expectation over $p$ and $D(\cdot\|\cdot)$ is the Kullback-Leibler (KL) divergence. EP vanishes in equilibrium, when trajectory statistics are indistinguishable under the forward and reverse distributions.
[…]
Our goal is to estimate $\sigma({{\bm{x}}})$ and $\Sigma$ from forward and reverse samples of observables $\bm{g}$. The overall irreversibility encoded in these observables is quantified by the KL divergence between the distributions of $\bm{g}$ under the forward and reverse processes, ${\Sigma}_{\bm{g}}^{\mathrm{DPI}}\vcentcolon=D(p_{G}\|\tilde{p}_{G})$, where $p_{G}(\bm{g}^{\prime})=\left\langle\delta_{\bm{g}^{\prime},\bm{g}({{\bm{x}}})}\right\rangle_{p}$ and $\tilde{p}_{G}(\bm{g}^{\prime})=\left\langle\delta_{\bm{g}^{\prime},\bm{g}({{\bm{x}}})}\right\rangle_{\tilde{p}}$. The data processing inequality (DPI) implies that ${\Sigma}_{\bm{g}}^{\mathrm{DPI}}\leq{\Sigma}$, with equality when $\bm{g}({{\bm{x}}})$ is an invertible function [36].
Estimating ${\Sigma}_{\bm{g}}^{\mathrm{DPI}}$ is very difficult for high-dimensional observables $\bm{g}$. Instead, we bound EP using a nonequilibrium generalization of the MaxEnt variational principle. Specifically, we choose the distribution that minimizes KL divergence to the reverse process while matching the forward-process expectations of the observables:
${\Sigma}_{\bm{g}}$ $\vcentcolon=\min_{q}D(q\|\tilde{p})\quad\text{subject to}\quad\left\langle\bm{g}\right\rangle_{q}=\left\langle\bm{g}\right\rangle_{p}.$ (2)
This quantity obeys the bounds $0\leq{\Sigma}_{\bm{g}}\leq{\Sigma}_{\bm{g}}^{\mathrm{DPI}}$, as shown in the Supplemental Material (SM) [37]. ${\Sigma}_{\bm{g}}$ quantifies the irreversibility captured by observable expectations, and it vanishes if and only if the expectation of $\bm{g}$ is the same under the forward and reverse processes. Large-deviations theory gives ${\Sigma}_{\bm{g}}$ a physical interpretation in terms of fluctuations: given $n$ sample trajectories from the reverse process, the probability that the empirical average of observables $\bm{g}$ is equal to $\left\langle\bm{g}\right\rangle_{p}$ scales as $\asymp e^{-n{\Sigma}_{\bm{g}}}$ [38].
Importantly, by exploiting a remarkable information-theoretic duality, we may calculate ${\Sigma}_{\bm{g}}$ without optimizing or inferring any probability distributions. As shown in SM [37], the optimization problem (2) has the dual formulation
[…]
Eq. (3) is an unconstrained convex optimization problem over Lagrangian multipliers $\bm{\theta}$} that enforce the $d$ expectation constraints on $\bm{g}$. The objective depends only on the expectations of $\bm{g}$ and $e^{{\bm{\theta}}^{\top}\bm{g}}$ under the forward and backward processes, which can be easily computed from empirical samples of $\bm{g}$. In the special case of antisymmetric observables and a steady-state system without odd variables, Eq. (3) can be estimated using only forward samples since $\ln\langle e^{{\bm{\theta}}^{\top}\bm{g}}\rangle_{\tilde{p}}=\ln\langle e^{-{\bm{\theta}}^{\top}\bm{g}}\rangle_{p}$. As we discuss in more detail below, a related (but different) variational bound on EP was proposed in Refs. [39, 40].
As an example, consider a system with 1000 binary spins measured at two timepoints, with $\bm{g}$ encoding two-point correlations $x_{i,t}x_{j,t^{\prime}}$. Eq. (3) is an unconstrained convex optimization problem over $1000^{2}$ variables, which can be solved using standard numerical techniques. Conversely, a naïve estimate of $\Sigma$} or ${\Sigma}_{\bm{g}}^{\mathrm{DPI}}$ requires inferrence of $2^{1000}$ probabilities.
So far, we have made no assumptions regarding multipartite structure. As we show in End Matter, in the presence of multipartite structure, where only a single observable can be active in a given trajectory, Eq. (3) can be split into smaller optimization problems, each involving a proportional fraction of variables and samples. This dramatically reduces the computation and memory required to solve Eq. (3) (see End Matter for an example).
Maximum likelihood EP decomposition.— As we show in the SM [37], the distribution $q^{*}$ that optimizes (2) belongs to an exponential family,
$q_{{\bm{\theta}}}({{\bm{x}}})=\tilde{p}({{\bm{x}}})e^{{{\bm{\theta}}}^{\top}\bm{g}({{\bm{x}}})-\ln\langle e^{{{\bm{\theta}}}^{\top}\bm{g}}\rangle_{\tilde{p}}}$. (4)
Specifically, $q^{*}=q_{{\bm{\theta}}^{*}}$ for optimal parameters $\bm{\theta}$$^{*}$ that assign maximum likelihood (ML) to forward trajectories:
${\bm{\theta}}^{*}=\operatorname*{arg\,max}_{{\bm{\theta}}\in\mathbb{R}^{d}}\,\big{\langle}\ln q_{{\bm{\theta}}}({{\bm{x}}})\big{\rangle}_{p}=\operatorname*{arg\,min}_{{\bm{\theta}}\in\mathbb{R}^{d}}\,D(p\|q_{{\bm{\theta}}}).$ (5)
This also leads to the information-geometric Pythagorean theorem [41], which decomposes EP into two nonnegative terms:
$\underbrace{D(p\|\tilde{p})}_{{\Sigma}}=$ $\underbrace{D(q^{*}\|\tilde{p})}_{{\Sigma}_{\bm{g}}}+\underbrace{D(p\|q^{*})}_{{\Sigma}^{\perp}_{\bm{g}}}.$ (6)
${\Sigma}^{\perp}_{\bm{g}}\vcentcolon={\Sigma}-{\Sigma}_{\bm{g}}=D(p\|q^{*})$ quantifies the EP not captured by ${\Sigma}_{\bm{g}}$, expressed via the ML problem ${\Sigma}^{\perp}_{\bm{g}}=\min_{{\bm{\theta}}}D(p\|q_{\bm{\theta}})$.
We may further decompose ${\Sigma}^{\perp}_{\bm{g}}$ by using that Eq. (4) implies that $q^{*}$ obeys $q^{*}_{X|G}({{\bm{x}}}|\bm{g}^{\prime})=\tilde{p}_{X|G}({{\bm{x}}}|\bm{g}^{\prime})$, where conditional probabilities are computed using Bayes’ rule, $\tilde{p}_{X|G}({{\bm{x}}}|\bm{g}^{\prime})=\tilde{p}({{\bm{x}}})\delta_{\bm{g}^{\prime},\bm{g}({{\bm{x}}})}/\tilde{p}_{G}(\bm{g}^{\prime})$. Using the chain rule of KL divergence, we may write ${\Sigma}^{\perp}_{\bm{g}}$ as a sum of two terms:
${\Sigma}^{\perp}_{\bm{g}}=\underbrace{D(p_{G}\|q^{*}_{G})}_{{\Sigma}_{\bm{g}}^{\mathrm{DPI}}-{\Sigma}_{\bm{g}}}+\underbrace{D(p_{X|G}\|\tilde{p}_{X|G})}_{{\Sigma}-{\Sigma}_{\bm{g}}^{\mathrm{DPI}}}\,.$ (7)
The first term is the additional EP that could be inferred by constraining all statistics (not just the mean) of $\bm{g}$ in Eq. (2). The second term is the EP that cannot be inferred from any statistics of $\bm{g}$.
[…]
Finally, our method gives an estimate of trajectory EP,
$\sigma_{{\bm{\theta}}^{*}}({{\bm{x}}})\vcentcolon=\ln\frac{q_{{\bm{\theta}}^{*}}({{\bm{x}}})}{\tilde{p}({{\bm{x}}})}={\bm{\theta}}^{*\top}\bm{g}({{\bm{x}}})-\ln\langle e^{{\bm{\theta}}^{*\top}\bm{g}}\rangle_{\tilde{p}}\,$. (8)
This estimator satisfies $\langle\sigma_{{\bm{\theta}}^{*}}\rangle_{p}={\Sigma}_{\bm{g}}$, and it may be shown to provide the best approximation of trajectory EP $\sigma({{\bm{x}}})$ within the exponential family $q_{\bm{\theta}}$. Specifically, if we use $\sigma_{{\bm{\theta}}}({{\bm{x}}})=\ln[q_{{\bm{\theta}}}({{\bm{x}}})/{\tilde{p}}({{\bm{x}}})]$ to indicate the approximation given by parameters $\bm{\theta}$} and $\delta\sigma_{{\bm{\theta}}}({{\bm{x}}})=\sigma({{\bm{x}}})-\sigma_{{\bm{\theta}}}({{\bm{x}}})$ the corresponding residual, then $\sigma_{{\bm{\theta}}^{*}}$ minimizes the expected error  [^1] Eq. (9) is derived by using $\delta\sigma_{{\bm{\theta}}}({{\bm{x}}})=\ln p({{\bm{x}}})/q_{{\bm{\theta}}}({{\bm{x}}})$ and $\langle e^{-\delta\sigma_{{\bm{\theta}}}}\rangle_{p}=1$.:
$D(p\|q_{{\bm{\theta}}})=\langle\delta\sigma_{{\bm{\theta}}}+e^{-\delta\sigma_{{\bm{\theta}}}}-1\rangle_{p}.$ (9)
Here, $\delta\sigma_{{\bm{\theta}}}+e^{-\delta\sigma_{{\bm{\theta}}}}-1\geq 0$ is an information-theoretic loss function that converges to squared error $(\delta\sigma_{{\bm{\theta}}})^{2}/2$ as $\delta\sigma_{{\bm{\theta}}}\to 0$.
[…]
$\sigma({{\bm{x}}})=\sigma_{{\bm{\theta}}^{*}}({{\bm{x}}})={\bm{\theta}}^{*\top}[\bm{g}({{\bm{x}}})-\bm{g}(\tilde{\bm{x}})]/2\,$, (10)
which simplifies further to ${\bm{\theta}}^{*\top}\bm{g}({{\bm{x}}})$ when $\bm{g}$ is antisymmetric.
Thermodynamic uncertainty relations (TURs).— Eq. (3) can be interpreted as a TUR that relates EP and fluctuations of trajectory observables [44]. Consider any trajectory observable that can be expressed as a linear combination $o({{\bm{x}}})={\bm{\theta}}^{\top}\bm{g}({{\bm{x}}})$ for some ${\bm{\theta}}\in\mathbb{R}^{d}$. The first term in (3) is the expectation of $o({{\bm{x}}})$, while the second term is the cumulant generating function (CGF) of $o({{\bm{x}}})$ under $\tilde{p}$. Thus, ${\Sigma}_{\bm{g}}$ bounds the fluctuation-discounted expectation of all such $o({{\bm{x}}})$, and it can be understood as a higher-order TUR [34] that constrains all cumulants of trajectory observables. This contrasts with quadratic TURs, which constrain only the mean and variance [45, 18, 46].
To relate our approach to quadratic TURs, we expand the CGF as $\ln\langle e^{{\bm{\theta}}^{\top}\bm{g}}\rangle_{\tilde{p}}\approx{\bm{\theta}}^{\top}\left\langle\bm{g}\right\rangle_{\tilde{p}}+{\bm{\theta}}^{\top}\mathsf{K}_{\tilde{p}}\,{\bm{\theta}}/2$, where $\mathsf{K}_{\tilde{p}}$ is the covariance matrix of $\bm{g}$ under $\tilde{p}$. Plugging into (3) gives the approximate optimizer $\hat{{\bm{\theta}}}$ as the solution to the linear system $\mathsf{K}_{\tilde{p}}\,\hat{{\bm{\theta}}}=\left\langle\bm{g}\right\rangle_{p}-\left\langle\bm{g}\right\rangle_{\tilde{p}}\equiv\left\langle\bm{g}\right\rangle_{p-\tilde{p}}$, giving the weaker bound
${\Sigma}_{\bm{g}}\geq\widehat{{\Sigma}}_{\bm{g}}$ $\vcentcolon=\left\langle\bm{g}\right\rangle_{p-\tilde{p}}^{\top}\mathsf{K}_{\tilde{p}}^{-1}\left\langle\bm{g}\right\rangle_{p}-\ln\left\langle e^{\left\langle\bm{g}\right\rangle_{p-\tilde{p}}^{\top}\mathsf{K}_{\tilde{p}}^{-1}\bm{g}}\right\rangle_{\tilde{p}}$. (11)
$\widehat{{\Sigma}}_{\bm{g}}$ can be shown to correspond to a single Newton-Raphson step for maximizing (3) starting from ${\bm{\theta}}=\bm{0}$ [37]. This bound has the advantage of not requiring optimization, only the solution of a linear system. The inequality ${\Sigma}_{\bm{g}}\geq\widehat{{\Sigma}}_{\bm{g}}$ is tight when $\bm{g}$ has Gaussian statistics.
We may also compare our results to existing TURs. We consider the special case of steady-state system without odd variables and antisymmetric $\bm{g}$. In this case, EP may be bounded as (see End Matter)
${\Sigma}\geq{\Sigma}_{\bm{g}}^{\textrm{TUR}}\vcentcolon=\ln\left(1+2\langle\bm{g}\rangle_{p}^{\top}\mathsf{K}_{p}^{-1}\langle\bm{g}\rangle_{p}\right)$ (12)
This bound approaches the usual quadratic form, ${\Sigma}_{\bm{g}}^{\textrm{TUR}}\simeq 2\langle\bm{g}\rangle_{p}^{\top}\mathsf{K}_{p}^{-1}\langle\bm{g}\rangle_{p}$ for sufficiently small $\langle\bm{g}\rangle_{p}^{\top}\mathsf{K}_{p}^{-1}\langle\bm{g}\rangle_{p}$. This regime holds in the short-time limit, since $\langle\bm{g}\rangle_{p}$ scales with the observation period.
The bounds (11)-(12) may require solving linear systems involving very large covariance matrices. For multipartite observables, it is often possible to split these bounds into sub-problems that involve smaller matrices (see End Matter).
[…]
where $v_{\bm{\phi}}({{\bm{x}}})$ is a nonlinear function parameterized by $\bm{\phi}\in\mathbb{R}^{k}$ (e.g., output of a neural network)  [47]. ${\Sigma}_{\bm{g}}$ in Eq. (3) is a special case of $\Sigma_{\bm{g}}^{\text{B}}$ where $v_{\bm{\phi}}({{\bm{x}}})$ contains all linear combinations of $\bm{g}({{\bm{x}}})$. The neural-network-based bound (13) has both practical and conceptual differences with respect to Eq. (3). For instance, it generally involves a difficult non-convex optimization problem, it does not provide a Pythagorean decomposition of EP, and it cannot be interpreted in terms of large deviations. Exploration of $\Sigma_{\bm{g}}^{\text{B}}$ for thermodynamic inference is left as an interesting future direction.
A related variational expression for thermodynamic inference was developed in Kim et al. [39] and Otsubo et al. [40]. In our notation, it is written as
$\Sigma_{\bm{g}}^{\text{KO}}\vcentcolon=\max_{{\bm{\theta}}}\left({\bm{\theta}}^{\top}\langle\bm{g}\rangle_{p}-\langle e^{{\bm{\theta}}^{\top}\bm{g}}\rangle_{\tilde{p}}+1\right)\,.$ (14)
It is possible to consider a neural-network-based version of $\Sigma_{\bm{g}}^{\text{KO}}$ by replacing ${\bm{\theta}}^{\top}\bm{g}({{\bm{x}}})$ with a parameterized function $v_{\bm{\phi}}({{\bm{x}}})$, as in Eq. (13), as explored in Refs. [39, 49].
Note that ${\Sigma}_{\bm{g}}\geq\Sigma_{\bm{g}}^{\text{KO}}$, since $\ln x\geq-x+1$ [47, 50], therefore ${\Sigma}_{\bm{g}}$ always provides a tighter bound on EP than $\Sigma_{\bm{g}}^{\text{KO}}$. The two bounds become equivalent in the limit $\langle e^{{\bm{\theta}}^{\top}\bm{g}}\rangle_{\tilde{p}}\to 1$, such as the short-time limit $T\to 0$ with antisymmetric observables when $\langle e^{{\bm{\theta}}^{\top}\bm{g}}\rangle_{\tilde{p}}=1+O(T)$ [40]. In general, however, the two bounds give different results, and in the End Matter, we provide a simple example where ${\Sigma}_{\bm{g}}$ is arbitrarily better than $\Sigma_{\bm{g}}^{\text{KO}}$. We note that outside the short-time limit, $\Sigma_{\bm{g}}^{\text{KO}}$ does not have a straightforward interpretation in terms of maximum likelihood inference or large-deviations statistics.
Lynn et al. [51, 52, 32] proposed a way to decompose and bound steady-state EP using a different information-theoretic optimization. Although originally focused on local EP in systems with multipartite dynamics, it can be generalized to arbitrary observables $\bm{g}$ and non-multipartite systems as
[…]
Although Eq. (15) resembles Eq. (2), it differs in that it simultaneously optimizes both arguments of the KL divergence, $q$ and $\tilde{q}$. As discussed in the End Matter, the optimal distribution in Eq. (15) is not in an exponential family, and in general it lacks full support. For this reason, Eq. (15) does not have a tractable dual expression analogous to our dual (3), and there is no straightforward way to scale ${\Sigma}_{\bm{g}}^{\textrm{L}}$ to large systems.
Other work has explored objective (15) in combination with other types of constraints, e.g., on waiting times [19, 14] and hidden Markovian structure [23, 13]. However, the resulting optimization typically does not scale to high-dimensional or even moderately-sized systems, in part because the considered constraints lead to non-convex optimization problems.
Our Pythagorean theorem (6) is related to previous decompositions of EP in interacting systems [53] and systems with nonconservative forces [54, 55, 35]. Moreover, the general form of Eq. (2) is a Maximum Entropy (MaxEnt) problem over trajectory distributions [56, 57, 58, 59], sometimes called “maximum caliber” (MaxCal) [60]. Related techniques have been used to infer models from which the entropy flow can be estimated, though without ensuring a lower bound on EP [61].
Our approach differs in several other ways from earlier work on MaxEnt and MaxCal. First, we minimize the KL divergence relative to an unknown prior distribution $\tilde{p}$, from which we typically only have samples. This makes standard approximation methods for large-scale MaxEnt problems, such as mean-field and Bethe approximations [62, 63, 26], not directly applicable. Second, we care not only about the parameters $\bm{\theta}$}, as is typical of MaxEnt “inverse problems”, but also about bounding EP by the quantity ${\Sigma}_{\bm{g}}$. Both of these issues are resolved by the dual formulation (3).
Other notable approaches to EP inference have employed compression algorithms [64, 65, 66] and deep learning of probability flows [67].
Examples: Nonequilibrium spin model and Neuropixels dataset.— We illustrate our method on two examples: a nonequilibrium kinetic Ising model [68, 69, 26] and in vivo spike data from the Neuropixels Visual Behavior repository [70]. For both examples, we consider a single discrete time step and binary variables, ${{\bm{x}}}=(\bm{x}_{0},\bm{x}_{1})$ where $x_{i,t}\in\{-1,+1\}$ for $i\in\{1,\dots,N\}$. We emphasize that our approach also works for continuous systems. In the SM [37], we illustrate it on a linear Langevin system.
To calculate ${\Sigma}_{\bm{g}}$, we sample from steady state using Monte Carlo. We then optimize Eq. (3) using gradient ascent with Barzilai–Borwein step sizes [71]. We perform early-stopping using held-out validation data, which avoids overfitting even in the far-from-equilibrium regime where many reverse transitions are not sampled. Reported EP estimates are computed on held-out test data. Details of the data generation, optimization, and analysis can be found in the SM [37] and our code repository [72].
The nonequilibrium spin model is specified by transition probabilities
$T(\bm{x}_{1}|\bm{x}_{0})=\frac{1}{N}\sum_{i}\Big{[}W_{i}(\bm{x}_{0})\delta_{\bm{x}_{0},\bm{x}_{1}^{[i]}}+(1-W_{i}(\bm{x}_{0}))\delta_{\bm{x}_{0},\bm{x}_{1}}\Big{]}$
where we introduced the spin-flip operator: $(\bm{x}^{[i]})_{i}=-x_{i}$ and $(\bm{x}^{[i]})_{j}=x_{j}$ for $j\neq i$. The flip probability for spin $i$ is
$W_{i}(\bm{x})$ $=\dfrac{\exp(-\beta x_{i}\sum_{j:j\neq i}w_{ij}x_{j})}{2\cosh(\beta\sum_{j:j\neq i}w_{ij}x_{j})}\,$, (16)
where $\beta$ is an inverse temperature and $w_{ij}$ are (typically asymmetric) coupling parameters. We consider the diluted version of the model [73, 74, 75] with $k$ average neighbors. Here $w_{ij}=c_{ij}z_{ij}/\sqrt{k}$, with binary connections $c_{ij}\sim\textrm{Bernoulli}(k/(N-1))$ and real-valued weights $z_{ij}\sim\mathcal{N}(0,1)$.
The steady-state EP in this model has a simple closed-form expression, ${\Sigma}=\beta\sum_{i\neq j}w_{ij}\langle{(x_{i,1}-x_{i,0})x_{j,0}}\rangle_{\rm st}$, where $\langle\cdot\rangle_{\rm st}$ indicates expectations under the stationary process $p$. This provides ground truth to evaluate our estimators.
Refer to caption(a)Refer to caption(b)
Figure 1: Disordered nonequilibrium spin model with 1000 spins. (a) Steady-state EP estimates for different inverse temperatures $\beta$. (b) Asymmetry of inferred parameters, shown against the true coupling asymmetries in the model for $\beta=2.5$ ($R^{2}=0.9831$). Estimates are based on $10^{9}$ state transitions sampled by Monte Carlo.
Our observables of interest are time-lagged correlations,
$g_{ij}({{\bm{x}}})=(x_{i,1}-x_{i,0})x_{j,0}\qquad\text{for all}\quad i,j\,$, (17)
with conjugate parameters $\theta_{ij}$. Because the system has multipartite observables, Eq. (3) can be decomposed into $N$ independent problems, each involving only those transitions where spin $i$ changes, improving computational performance for large systems (see SM).
Fig. 1(a) shows the actual and inferred EP at different $\beta$ for $N=1000$, $k=6$. EP increases with $\beta$, and all estimators (${\Sigma}_{\bm{g}}^{\textrm{TUR}}$, $\widehat{{\Sigma}}_{\bm{g}}$, ${\Sigma}_{\bm{g}}$) agree in the near-equilibrium regime of small $\beta$. Importantly, ${\Sigma}_{\bm{g}}$ provides a tight bound on EP even in the far-from-equilibrium regime of large $\beta$. The gap between ${\Sigma}_{\bm{g}}$ and $\widehat{{\Sigma}}_{\bm{g}}$ indicates the onset of highly non-Gaussian statistics from $\beta\approx 2$.
As discussed above, ${\Sigma}_{\bm{g}}\approx{\Sigma}$ implies that the trajectory EP $\sigma({{\bm{x}}})$ is closely approximated by $\sigma_{{\bm{\theta}}^{*}}$ from Eq. (8). As shown in the SM, this allows us to use the optimal parameters $\bm{\theta}$$^{*}$ to infer the asymmetry of the coupling constants, $\theta_{ij}^{*}-\theta_{ji}^{*}\approx\beta(w_{ij}-w_{ji})$. This relation is verified in Fig. 1(b).
In the SM, we also consider other coupling matrices for which the model exhibits a nonequilibrium phase transition [26]. We show that our method accurately infers EP in this regime, a difficult task for existing approximations [76].
(a)Refer to captionRefer to caption(b)
Figure 2: (a) Estimated EP per expected number of spikes $R$, in the Neuropixels Visual Behavior dataset for three conditions. Error bars indicate standard error of the mean. (b) A sample of inferred coupling coefficients $\theta^{*}_{ij}$ grouped by visual area. Here we select 200 neurons with highest firing rate from an active trial. To improve visualization, lower triangle shows $\theta_{ji}^{*}\equiv-\theta_{ij}^{*}$ for $i<j$.
Next, as an example application to biological data, we estimate EP in the Visual Behavior Neuropixels dataset [70], which contains spike-train recordings from 81 mice in 103 sessions. In this example, EP is understood as a statistical measure of temporal irreversibility, rather than energetic dissipation [33]. The dataset includes spiking activity from multiple brain regions, including visual cortical areas (VISp, VISl, VISal, VISrl, VISam and VISpm) and subcortical structures.
We analyze data from visual areas during three conditions: active behavior (visual change detection task), passive replay (same stimuli but without task engagement), and Gabor (receptive field characterization with Gabor stimuli and full-field flashes). We discretize spike trains into temporal bins of length $10$ ms and verify that most bins contain no more than one spike. In this dataset, neurons can update in a parallel (non-multipartite) manner.
Our observables are defined as
$g_{ij}({{\bm{x}}})=x_{i,1}\,x_{j,0}-x_{i,0}\,x_{j,1}\qquad\text{for all}\quad i<j\,$, (18)
which naturally represent time-lagged correlations in non-multipartite dynamics. They capture the antisymmetric part of (17) and give more robust numerical results on this dataset.
We estimate EP across different conditions, recording sessions, and system sizes $N$. We randomly select 10 subsets of neurons and estimate ${\Sigma}_{\bm{g}}$ for each subset. For improved comparison, we normalize EP in each condition by the expected number of spikes per bin, $R=\sum_{i}\left(1+\langle x_{i,0}\rangle_{\rm st}\right)/2$.
Fig. 2(a) shows that EP grows superlinearly with size (note that $R\propto N$). It also shows that the active condition is associated with the largest normalized EP. Fig. 2(b) illustrates the matrix of inferred parameters $\theta_{ij}^{*}$ for 200 neurons from an active trial, grouped by visual brain area. This matrix specifies a network of functional connectivity [77], showing how interactions between individual neurons contribute to temporal irreversibility of in vivo brain dynamics. This network reveals a clustered organization aligned with anatomical brain regions.
Author Contributions
A.K. and M.A. contributed equally to this work.
Acknowledgements
Acknowledgements.
M.A. is partly supported by John Templeton Foundation (grant 62828) and Grant PID2023-146869NA-I00 funded by MICIU/AEI/10.13039/501100011033 and cofunded by the European Union, and supported by the Basque Government through the BERC 2022-2025 program and by the Spanish State Research Agency through BCAM Severo Ochoa excellence accreditation CEX2021-01142-S funded by MICIU/AEI/10.13039/501100011033. S. I. is supported by JSPS KAKENHI Grants No. 22H01141, No. 23H00467, and No. 24H00834, JST ERATO Grant No. JPMJER2302, and UTEC-UTokyo FSI Research Grant Program. A.K. is partly supported by John Templeton Foundation (grant 62828) and by the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie Grant Agreement No. 101068029.
I Data Availability Statement
The code used in this study is available at a GitHub repository [72].